OpenStudy (anonymous):

Solve for y y+4/2y + y-2/ 3 = 3y^2+10/ 6y

OpenStudy (mathteacher1729):

How far did you get before getting stuck? Also, do you know how to solve this graphically?

OpenStudy (anonymous):

Using parenetheses would be helpful

OpenStudy (anonymous):

(y+2)/(2y) + (y-2)/(3) = (3y^2+10)/(6y)

OpenStudy (anonymous):

thank you

myininaya (myininaya):

\[\frac{y+2}{2y}+\frac{y-2}{3}=\frac{3y^2+10}{6y}\] \[6y*\frac{y+2}{2y}+6y*\frac{y-2}{3}=6y*\frac{3y^2+10}{6y}\]

OpenStudy (anonymous):

6y(y+2)(y-1)

OpenStudy (mathteacher1729):

If you go to http://www.wolframalpha.com and type solve (y+2)/(2y) + (y-2)/(3) = (3y^2+10)/(6y) and then click "Show steps" you can get a detailed breakdown of the solution. It gives a complex answer, though, so I'm wondering if you copied the problem correctly?

OpenStudy (anonymous):

lol

myininaya (myininaya):

3(y+2)+2y(y-2)=3y^2+10

myininaya (myininaya):

3y+6+2y^2-4y=3y^2+10 3y^2+10-3y-6-2y^2+4y=0 y^2+y+4=0 you can solve the rest of this using the quad formula

OpenStudy (anonymous):

He changed the problem

OpenStudy (anonymous):

The first term was y+4 in teh denom, then it became y+2

OpenStudy (anonymous):

:/

myininaya (myininaya):

good eye so we have two problems here and i did the second one not the first one

OpenStudy (anonymous):

I did the first one, lol

OpenStudy (anonymous):

It's the original one.

OpenStudy (anonymous):

That website helped though.

OpenStudy (anonymous):

If it's the original one, then mine is right.

myininaya (myininaya):

mine is right lol